The probability that George goes swimming on any day is 1/3. Use an approximation to calculate the probability that in 270 days George goes swimming at least 100 times.
- The probability that George goes swimming on any day is 1/3. Use an approximation to calculate the probability that in 270 days George goes swimming at least 100 times.[5]
Mark scheme
Marking points from the official mark scheme (May/June 2017). The mark or mark type is shown in brackets.
- [B1] np = 270 × 1/3 = 90, npq = 270 × 1/3 × 2/3 = 60 Correct unsimplified np and npq, SOI
- [M1] P(x > 100) = P(z > (99.5-90)/sqrt(60)) = P(z > 1.2264) Standardising using 100 need sq rt
- [M1] Continuity correction, 99.5 or 100.5 used
- [M1] = 1-0.8899 Correct area 1 – Φ implied by final prob. < 0.5
- [A1] = 0.110
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →